When an AC source is connected across a 10.0 Ω resistor, the output voltage is given by Δv = (140 V)sin(50πt). Determine the following quantities.

(a) maximum voltage
V

(b) rms voltage
V

(c) rms current
A

(d) peak current
A

(e) Find the current when t = 0.0045 s.
A
Answer #1

A)

Maximum voltage, V = 140 V

B)

RMS_voltage = 140/sqrt(2) = 99 V

C)

RMS_current = 99/10 = 9.9 A

D)

Maximum current, I = 140/10 = 14 A

E)

I = 14 sin(50 x pi x 0.0045) = 9.09 A

Comment in case any doubt please rate my answer....